### v1

# A Robust Autoencoder Ensemble-Based Approach for Anomaly Detection in Text

Abstract
In this work, a robust autoencoder ensemble-based approach designed to address anomaly detection in text corpora is introduced. Each autoencoder within the ensemble incorporates a local robust subspace recovery projection of the original data in its encoding embedding, leveraging the geometric properties of the k-nearest neighbors to optimize subspace recovery and identify anomalous patterns in textual data. The evaluation of such an approach needs an experimental setting dedicated to the context of textual anomaly detection. Thus, beforehand, a comprehensive real-world taxonomy is introduced to distinguish between independent anomalies and contextual anomalies. Such a study to identify clearly the kinds of anomalies appearing in a textual context aims at addressing a critical gap in the existing literature. Then, extensive experiments on classical text corpora have been conducted and their results are presented that highlights the efficiency, both in robustness and in performance, of the robust autoencoder ensemble-based approach when detecting both independent and contextual anomalies. Diverse range of tasks, including classification, sentiment analysis, and spam detection, across eight different corpora, have been studied in these experiments.
A Robust Autoencoder Ensemble-Based Approach for Anomaly Detection in Text
Jeremie Pantin  and Christophe Marsala Sorbonne Université - LIP6 Paris - France Firstname.Lastname@lip6.fr
1 Introduction
Anomaly detection Chandola et al. (2009); Ruff et al. (2021) and outlier detection Hawkins (1980); Hodge and Austin (2004); Zhang (2013); Aggarwal (2017a) are tasks aiming to estimate whether an observation is normal or not. Both of these tasks are referring to similar characteristics, and the terms anomaly and outlier are interchangeable in this paper. Several works have been devoted to study and categorise the different characteristics of an anomaly, and three principal kinds of anomalies can be noted: independent, contextual and collective. Defining a taxonomy for anomalies is crucial while the task to find anomalies lacks dedicated datasets. Such issue often leads practitioners to introduce an artificial contamination to produce anomalies into datasets. With the emergence of new machine learning methods and the availability of many datasets and corpora, anomaly detection can be addressed through many approaches Beggel et al. (2020); He et al. (2021); Goodge et al. (2023), however, in most cases, models are based on one-class classification (OCC) Khan and Madden (2014); Ruff et al. (2018); Sohn et al. (2021).
Figure 1: Presentation of the studied problem with three documents topics: medical, culture and politic. Under each topic we represent a textual document with colored rectangles. Gray and green are inliers and red ones are anomalies. The detailed documents are the abstract of the news articles taken from sources like Reuters, New York times, BBC, … The first scenario is the apparition of a culture-related document in a medical feed, and the second scenario is a political document in the culture feed.
Performing anomaly detection on textual data is less common than on many other types of data (images, time series and medical) but it comes with several useful applications that helps discerning wrong web content, hateful messages, spam or also errors in news feed. One of the main difficulties in this domain is the reproduction of experimental protocols and results from the literature Bejan et al. (2023). Indeed, there is a great difference between tackling independent outliers and contextual outliers Mahapatra et al. (2012); Fouché et al. (2020); Bejan et al. (2023) using semantic in text. For the former, anomaly classifier needs to differentiate two kinds of documents that come from unrelated topics (eg. sports and computer) but for the latter, they needs to detect a topic contaminated with another "sibling" (see Fig. 1 for examples of this difference). It appears that most of the recent works contaminate corpora without addressing the problem of which kind of anomaly/outlier is added in this process Manevitz and Yousef (2001); Kannan et al. (2017); Ruff et al. (2019a); Lai et al. (2020).
In this paper, based on recent literature, we address the problem of taxonomy after defining two kinds of anomaly: independent and contextual. Those anomalies are dependent of any semantic hierarchy and we propose in the experiments section to detail furthermore how they are benchmarked. For tackling both kind of anomalies, we propose a novel ensemble-based approach built with a robust autoencoder.
This paper is composed as follows. In Section 2, recent works are presented. In Section 3, our approach to build an Ensemble robust subspace local recovery encoder is introduced with a description of its properties. Several experiments are presented in Section 4. Finally, Section 6 concludes and presents some future work.
2 Related works
Recent advances in word embedding with language models like BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019), and large language models such as Llama Touvron et al. (2023) and OPT Zhang et al. (2022) have shown promising characteristics for anomaly detection. However, only few methods of the literature propose their use Manolache et al. (2021); Ruff et al. (2019a). Other methods like One-Class Support Vector Machine (OCSVM) Schölkopf et al. (2001) and Textual Outlier using Nonnegative Matrix Factorization (TONMF) Kannan et al. (2017) rely on a TF-IDF representation. Surprinsinly, recent methods are not using outlier ensemble methods Aggarwal and Sathe (2015); Zhao et al. (2019a); Zimek et al. (2014) for performing anomaly/outlier detection with text data. Their usage in both Mixture of Experts Du et al. (2022) and ensemble approach Briskilal and Subalalitha (2022) in various NLP tasks highlights that anomaly detection in text data misses important approaches.
AutoEncoders (AE) have been widely used for anomaly/outlier detection with high-dimensional data Chen et al. (2017); Kieu et al. (2019) and are also successful with other kind of data An and Cho (2015); Chen et al. (2018); Lai et al. (2020); Zhou and Paffenroth (2017); Beggel et al. (2020). The risk of using autoencoders with language models leads to the apparition of degenerate solution in the learning step. Robust properties are needed in such scenario for handling the manifold collapse phenomenon Wang and Wang (2019). Such issue can be mitigated using robust projections in the learned space of the autoencoder (eigenfunction decomposition Bengio et al. (2004), …). The Robust Subspace Recovery Lerman and Maunu (2018); Rahmani and Atia (2017) is a robust manifold learning technique that map inlier distributions in a subspace where anomalies/outliers are at the edge. An autoencoder based on such an approach Lai et al. (2020) encodes observation from the input space in a robust subspace. One problem of such autoencoder is the difficulty to address scenarios with features that present similar range values (problem of entanglement). This entanglement is especially more difficult with contextual anomalies where documents shares similar semantic properties. Such problem can be tackled using a local projection from the nearest neighbours in the latent representation space. Several manifold learning approaches like locally linear embedding (LLE) Roweis and Saul (2000), neighbourhood components analysis (NCA) Goldberger et al. (2004) or soft nearest neighbours (SNN) Frosst et al. (2019) can help to mitigate it.
In this paper we introduce an ensemble-based approach with randomly connected autoencoders. We also present a new autoencoder, robust subspace local recovery autoencoder (RLAE), that performs robust subspace representation with local neighbouring recovery. In this approach, called Ensemble robust subspace Local recovery Autoencoder (ErLA), a RLAE ensemble, it can be seen as an ensemble of several subspace that aims to find normal data with different manifolds. It is based on the hypothesis that one distribution is highly contaminated, and that inliers lies in a low-dimensional subspace.
3 ErLA: Ensemble robust subspace Local recovery Autoencoder
This section presents our approach ErLA and a description of its properties. While robust subspace recovery autoencoders have successfully tackled anomaly detection in text corpora, they lack locality and geometry awareness for mitigating disentanglement and manifold collapse in transformer-based language models. Thus, we introduce RLAE which integrates locality in the latent representation through a local neighbouring method.
The section is structured with a presentation of the randomly connected autoencoders, followed by a presentation of RSR loss. We then introduce the locally linear embedding loss term of ErLA before presenting its ensemble method. Finally, we present the representation of text.
3.1 Randomly Connected One-Class Autoencoder
Instead of using fully connected autoencoders, we propose to use randomly connected autoencoders. In the case of RSRAE, it is a novel approach and allow us to build ensemble autoencoders with different base detectors.
Let $\mathrm{X}$ be a dataset of $\mathrm{N}$ instances such as $\mathrm{X}=\{\mathrm{x_{1}},...,\mathrm{x_{N}}\}$. Each instance has $\mathrm{D}$ dimensions which correspond to its attributes: $\mathrm{x_{i}}=\{x_{1},...,x_{D}\}$. An Autoencoder is a neural networks in which the encoder $\mathcal{E}$ maps an instance $\mathrm{x_{i}}$ in a latent representation noted $\mathrm{z}_{i}=\mathcal{E}(\mathrm{x}_{i})\in\mathbb{R}^{e}$ of dimension $e$. The RSR layer is a linear transformation $\mathbf{A}\in\mathbb{R}^{d\times e}$ that reduces the dimension to $d$. We denote $\mathrm{\hat{z}_{i}}$ the representation of $\mathrm{z_{i}}$ through the RSR layer, such as $\mathrm{\hat{z}_{i}}=\mathbf{A}\mathrm{z_{i}}\in\mathbb{R}^{d}$. The decoder $\mathcal{D}$ maps $\mathrm{\hat{z}_{i}}$ to $\mathrm{\hat{x}_{i}}$ in the original space $D$. The matrix $\mathbf{A}$ and the parameters of $\mathcal{E}$ and $\mathcal{D}$ are obtained with the minimization of a loss function.
Similarly to Chen et al. (2017) we introduce autoencoders with random connection such as we increase the variance of our model. In the autoencoders ensemble each autoencoder has a random probability of having several of its connections to be cut. Thus, we set the probability of disconnection with a random rate in $[0.2,0.5]$.
3.2 Robust Subspace Recovery Layer
We present the RSR AutoEncoder that aims to robustly and nonlinearly reduce the dimension of the original data Lerman and Maunu (2018). The RSR layer maps the inliers around their original locations and the outliers far from their original locations. The loss function minimises the sum of the autoencoder loss function noted $L_{AE}$ with the RSR loss function noted $L_{RSR}$.
$L^{p}_{AE}(\mathcal{E},\mathbf{A},\mathcal{D})=\sum^{N}_{i=1}{||\mathrm{x_{i}}-\mathrm{\hat{x}_{i}}||^{p}_{2}}$ (1)
which is the $l_{2,p}-norm$ based loss function for $p>0$.
For performing the subspace recovery, we denote two terms that have different roles in the minimisation process. The first term enforces the RSR layer to be robust (PCA estimation) and the second enforces the projection to be orthogonal:
$L^{q}_{RSR}(\mathbf{A})=$ $\lambda_{1}\sum^{N}_{i=1}{||\mathrm{z_{i}}-\mathbf{A}^{\mathrm{T}}\mathrm{\hat{z}_{i}}||^{q}_{2}}$
$+\lambda_{2}\sum^{N}_{i=1}{||\mathbf{AA}^{\top}-\mathbf{I}_{d}||^{q}_{f}}$ (2)
with $\mathbf{A}^{\top}$ the transpose of $\mathbf{A}$, $\mathbf{I}_{d}$ the $d\times d$ matrix and $||\cdot||_{f}$ the Frobenius norm. $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters and $q=1$ is corresponding to the optimal $l_{p,q}$ norm Maunu et al. (2019). If we simplify Equation 2 we have:
$L_{RSRAE}(\mathcal{E},\mathbf{A},\mathcal{D})=$ $\lambda_{1}L^{1}_{AE}(\mathcal{E},\mathbf{A},\mathcal{D})$
$+\lambda_{2}L^{1}_{RSR}(\mathbf{A})$ (3)
3.3 Robust local embedding
Locally Linear Embedding (LLE) Roweis and Saul (2000); Chen and Liu (2011) is a popular dimensionality reduction technique that aims to preserve the local geometry of the data in a lower-dimensional subspace. It is based on the assumption that data points in a local neighbourhood can be linearly represented by their neighbouring data points. The LNE term in the loss function encourages the autoencoder to learn representations that preserve the relationships between data points in their local neighbourhoods. By doing so, it helps to project the Euclidean distance with its neighbours in the learned subspace. Based on Equation 2, the reconstruction loss function of RSRAE enforces robustness with $L^{1}_{AE}$ and the orthogonality with $L^{1}_{RSR}$. Because the learned representation of the encoder is compressed in a $e$ dimension space, the locality of the subspace is not handled.
For tackling this problem, we propose to introduce a third term to $L_{RLAE}$ based on a local neighbours embedding (LNE). Given a set of data points $\{\mathrm{x_{i}}\}_{i=1}^{N}$ in the input space, the goal of LNE is to find a lower-dimensional representation $\{\mathrm{z_{i}}\}_{i=1}^{N}$ in the output space (the subspace learned by the autoencoder) such that the local relationships between data points are preserved. We note:
$L_{LNE}(\mathbf{A})=\sum_{i=1}^{N}{\left\lVert\mathrm{x_{i}}-\sum_{j\in\mathcal{N}_{i}}w_{ij}\mathrm{x_{j}}\right\rVert_{2}^{2}}$ (4)
where $\mathcal{N}_{i}$ represents the set of indices of the k-nearest neighbours of $\mathrm{x_{i}}$ (excluding $\mathrm{x_{i}}$ itself) and $w_{ij}$ are the weights assigned to the neighbouring observations $\mathrm{x_{j}}$ in the reconstruction of $\mathrm{x_{i}}$. The weights $w_{ij}$ can be computed using the least squares method to minimize the reconstruction error: $\min_{\mathbf{w_{i}}}\left\lVert\mathrm{x_{i}}-\sum_{j\in\mathcal{N}_{i}}w_{ij}\mathrm{x_{j}}\right\rVert_{2}^{2}$ subject to the constraint $\sum_{j\in\mathcal{N}_{i}}w_{ij}=1$.
The LNE term encourages the autoencoder to find a representation for each data point as a combination of its k-nearest neighbours in the input space. With minimisation of the LNE term in the loss function, the autoencoder learns to preserve the local linear relationships, which ultimately helps to project the Euclidean distance with its neighbours. The reconstruction errors of LNE is measured by the cost function:
$L_{LNE}(\mathbf{A})=\sum^{N}_{i=1}{\sum_{j\in\mathcal{N}_{i}}w_{ij}\lVert\textbf{A}\mathrm{x_{i}}-\mathbf{A}\mathrm{x_{j}}\rVert_{2}^{2}}$ (5)
The weight $w_{j}$ assigned to the neighbour $\mathrm{x_{ij}}$ in the local reconstruction of $\mathrm{x_{i}}$ is determined based on its corresponding distances. The inclusion of the LNE term in the loss function encourages the autoencoder to preserve the local geometric structure of the data in the learned subspace.
Finally, the RLAE cost function is measured as follows:
$L_{RLAE}(\mathcal{E},\mathbf{A},\mathcal{D})$ $=L_{RSRAE}(\mathcal{E},\mathbf{A},\mathcal{D})$
[…]
The parameter $\lambda_{3}$ controls the influence of the LNE term on the overall loss. Because it controls the influence of locality of the manifold, the term is preferred to be low for avoiding degenerate results.
3.4 Ensemble Learning
The main idea behind ensemble methods is that a combination of several models, also called base detectors, and their outputs is more robust than usage of a single model. Such robustness can be observed against the bias-variance tradeoff and also for tackling the issue of overfitting.
Although the possibility to combine multiple base detectors is intuitive, the design of such approaches needs special attention regarding normalisation of outputs. In ErLA, we use the reconstruction error of each autoencoders and then we normalise each base detector scores through the standard deviation of one unit. We then take the median value for each observation.
3.5 Text Representation
Recent works Ruff et al. (2019a); Manolache et al. (2021) have recorded their results on BERT-based language models. Thus, we use RoBERTa Liu et al. (2019) for text representation. The ErLA model is not based on the self-attention mechanism, such as for Ruff et al. (2019a); Manolache et al. (2021), and we use the implementation of Reimers and Gurevych (2019).
4 Experiments
In this section we present several experiments on both independent anomalies and contextual anomalies. We present all corpora and how we perform this study. The section is structured as follows: i) presentation of the experimental protocol for tackling independent and contextual anomalies; ii) presentation of all corpora and their principal characteristics; iii) complete description of experiment settings; iv) results of conducted benchmarks.
Algorithm 1 TAC: Textual Anomaly Contamination
Inlier topic $\zeta$, corpus $\mathrm{X}$, split size $l$, contamination rate $\mathrm{\nu}$
$0<l\leq N$
$c\leftarrow l\nu$
$i\leftarrow 0$
Initialize empty matrix $\mathrm{Z}$
$\mathcal{A}\leftarrow\{\mathrm{x}_{j}\times\mathrm{y}_{j}\in\mathrm{X}\times\mathrm{Y}|\forall j\in[0,N],\mathrm{y}_{j}\neq\zeta\}$
$\mathrm{X}_{\zeta}\leftarrow\{\mathrm{X}\backslash\mathcal{A}\}$triangleright Inlier Matrix
while $|\mathrm{Z}|<c$ do
     if $\mbox{Parent}(\mathrm{y}_{i})\neq\mbox{Parent}(\zeta)\mbox{}$ then
         $\mbox{Append}(\mathrm{x}_{i},\mathrm{y}_{i})$ to $\mathrm{Z}$
     end if
     $i\leftarrow i+1$
end while
Fill $\mathrm{Z}$ with $\mathrm{X}_{\zeta}$ until $|\mathrm{Z}|=l$
return $\mbox{Shuffle}(\mathrm{Z})$
4.1 Experimental protocol
A large number of contributions have settled anomaly/outlier taxonomies Hawkins (1980); Hodge and Austin (2004); Zhang (2013); Aggarwal (2017b); Ruff et al. (2019a) several types of outliers have been proposed in the literature: Point anomaly, Conditional/Contextual anomaly and Collective/Group anomaly. A similar taxonomy can be applied to textual data. Consequently, various types of outliers frequently coexist within the documents of a given corpus. The definition of a topic can be assimilated to the subject matter that a document addresses. Depending on the document type, there may be multiple subtopics within a broader category (e.g., a sports topic that encompasses football and tennis). Thus, accounting for this hierarchical structure introduces a form of contextual anomaly. These contextual outliers may appear unremarkable in isolation but are considered outliers when associated with a small subset of the corpus.
Collective anomaly poses challenges in terms of formalisation due to the contextual nature of textual data. To illustrate this, consider a legal document mentioning a football player, which would be anomalous if incorrectly appearing in a sports-related corpus. Point outliers represent observations that lack any meaningful relationship with other topics. Specifically, anomaly topics and inlier topics have different hierarchical parents within the category structure. Let a labelled document of a corpus $(\mathrm{x},y)\in\mathrm{X}\times\mathcal{Y}$ and $\zeta$ be the inlier category, and its corresponding subset $\mathrm{X}_{\zeta}\subseteq\mathrm{X}$. We define $\mathcal{A}$ the subset of all anomalies such as $\mathcal{A}\subset\mathrm{X}$. We have:
$\mathcal{A}=\mathrm{X}\backslash\mathrm{X}_{\zeta}$ (7)
Regarding $\mathcal{A}$, we can make the distinction with two different constraints. We note $P(y)$ the direct parent of $y$ in a given hierarchy. First, an observation $\mathrm{x_{i}}$ is considered to be an anomaly if its parent topic is different of inlier parent topics such as:
$\mathcal{A}_{p}(\zeta)=\{\mbox{P}(\zeta)\neq\mbox{P}(\mathrm{y})|(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y}\}$ (8)
The second constraint corresponds to documents that do not lie in $\mathrm{X}_{\zeta}$ but share the same parent topic as $\zeta$. These observations are identified as another kind of anomaly: contextual anomaly. We write:
$\mathcal{A}_{c}(\zeta)=\{\mbox{P}(\zeta)=\mbox{P}(\mathrm{y})|(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y},\mathcal{A}\backslash\mathcal{A}_{p}\}$ (9)
Dataset Topics Hierarchy
20 Newsgroups 20 Yes
DBpedia 14 14 Yes
Reuters-21578 90 No
Web of Science 134 Yes
Enron 2 No
SMS Spam 2 No
IMDB 2 No
SST2 2 No
Table 1: Presentation of datasets from the literature on outlier detection and the inherent tasks. We describe these corpora by indicating the existence of a topic hierarchy in the labels of the original corpus.
4.2 Data
Although there are dedicated datasets for outlier detection, such as ODDS or UCI, they mainly provide multidimensional data, time series and computer vision data. Applications such as spam detection and text classification have a rich set of corpora available. Recent work Lai et al. (2020); Ruff et al. (2019a); Kannan et al. (2017); Mahapatra et al. (2012) uses classification datasets such as Reuters-21578111 http://www.daviddlewis.com/resources/testcollections/reuters21578/ and 20 Newsgroups222 http://qwone.com/~jason/20Newsgroups/ with a dedicated preparation in order to compare their approaches.
We use the corpus presented in the section and for each available category, we apply the preparation of independent outliers and contextual outliers with TAC (Algorithm 1). To be fair with each method and each dataset, we set the size of the preparation subset to $350$ and the results are averaged over $10$ of runs. The data is pre-processed by removing lower case and stop words. The train part of each corpus is used for training and evaluation. The TF-IDF model is applied to the whole corpus and only tokens that appear at least three times are kept in the vocabulary. In a first step, we set $\nu=0.10$.
Independent anomalies ($\nu=0.1$)
Model 20New. Reuters WOS DBpedia Enron SMS Sp. IMDB SST2
OCSVM 0.884 0.851 0.876 0.962 0.687 0.683 0.539 0.575
OC-AE 0.630 0.721 0.655 0.599 0.484 0.683 0.517 0.499
RandAE 0.622 0.689 0.636 0.575 0.466 0.687 0.503 0.486
RSRAE 0.669 0.768 0.731 0.634 0.518 0.642 0.540 0.577
CVDD 0.781 0.895 0.948 0.958 0.629 0.717 0.513 0.562
RLEA 0.792 0.829 0.843 0.840 0.611 0.657 0.521 0.542
ErLA 0.907 0.926 0.939 0.983 0.713 0.813 0.560 0.594
Contextual anomalies ($\nu=0.1$)
Model 20New. Reuters WOS DBpedia
AP AUC AP AUC AP AUC AP AUC
OCSVM 0.220 0.681 0.471 0.755 0.599 0.889 0.566 0.882
OC-AE 0.175 0.592 0.351 0.683 0.162 0.577 0.334 0.714
RandAE 0.164 0.579 0.337 0.662 0.147 0.544 0.319 0.694
RSRAE 0.177 0.584 0.340 0.729 0.214 0.623 0.260 0.660
CVDD 0.131 0.613 0.481 0.824 0.623 0.916 0.646 0.919
RLEA 0.201 0.642 0.420 0.757 0.334 0.738 0.368 0.747
ErLA 0.325 0.718 0.609 0.880 0.687 0.921 0.840 0.951
Table 2: Results of state of the art models for independent and contextual anomalies with the contamination rate $\nu=0.10$. Area under ROC (AUC) is the default evaluation metric for independent contamination, and Average Precision (AP) is also added for contextual contamination. The experimental study is performed on Distill RoBERTA. Each result is performed on test split prepared through Algorithm 1.
20 Newsgroups.
For 20 Newsgroups we separate the subtopics into seven main topics: computer, forsale, motors, politics, religion, science, sports. We do not count the topic "forsale" for contextual outliers because it has no subtopics.
Reuters-21578.
The Reuters-21578 corpus contains documents associated with several topics. We delete all these documents to keep only those associated with a single topic. We reorganise the topics in order to obtain a hierarchy, based on the work of Toutanova et al. (2001). Thus, four parent themes are created: commodities, finance, metals and energy. We apply TAC to the eight topics that have the greatest number of training documents.
DBpedia 14.
For DBpedia 14 we create the topic hierarchy based on the ontology provided333mappings.dbpedia.org/server/ontology/classes/ and has six parent topics.
Web Of Science.
Web of Science is often used as a reference for hierarchical classification and provides three levels of topic hierarchy. The third level topics are distributed among the corresponding first level parents. Thus, seven parent topics are present and for child topics that are associated with more than one parent, we keep the largest set of children and delete the others.
(a) DBpedia 14
(b) 20 Newsgroups
(c) Web of Science
(d) Reuters-21578
Figure 2: Results of our experimental study with $\nu=0.1$, split size of $350$ and number of base detector of $25$. The performance metric is AUROC (AC) and the text representation is RoBERTA.
(a) DBpedia 14
(b) 20 Newsgroups
(c) Web of Science
(d) Reuters-21578
Figure 3: Boxplots of results of our experimental study with $\nu=0.1$, split size of $350$ and number of base detector of $25$. The performance metric is AUPRC (AP) and the text representation is RoBERTA.
4.3 Setup
We use TAC for preparing contextual contamination on each candidate inliers possible with $\nu=0.1$ (TAC1) and a split size of min$(|D|,1000)$. All results are performed on AUROC and AUPRC reference works from the previous section. We integrate results of one-class autoencoder (OC-AE) and we also benchmark results on a randomly connected autoencoder ensemble (RandAE) Chen et al. (2017). The architecture is similar to Chen et al. (2017) and the autoencoders are following their settings. We also keep the number of runs for each corpus and each split contamination to $10$.
Implementation
For ErLA and RandAE we setup similarly than with the autoencoder and we set the number of base predictors to $20$. We provide the code of our approach [^4] https://github.com/jrmip/ErLA/ using the PyOD base implementation Zhao et al. (2019b). We propose an implementation based on the BaseDetector of PyOD so each of the compared model can be accordingly tested. Randomly connected neural networks can be difficult to implment. Thus, we have implemented three version of our randomly connected autoencoder: one with setting the prune connections to 0 before each backpropagation, another one with the usage of PyTorch built-in prune module and finally a version with Torch-Pruning Fang et al. (2023). Based on our experiments, we use the built-in Pytorch module for reproducibility. Additionally, we also set hyperparameters $\lambda_{1}=0.1$, $\lambda_{2}=0.1$ and $\lambda_{3}=0.05$. For avoiding manifold collapse problem and degenerates solutions, we advise that $\lambda_{3}<\lambda_{1}$. We set the latent dimension of the LNE layer to $32$. On the other hand, we set the epoch number to $50$ and random connection probability between $[0.2,0.5]$.
(a) DBpedia 14
(b) 20 Newsgroups
(c) Web of Science
(d) Reuters-21578
Figure 4: Number of base detectors considering AUC performances on contextual contamination.
(a) DBpedia 14
(b) 20 Newsgroups
(c) Web of Science
(d) Reuters-21578
Figure 5: Number of k-nearest neighbours for one RLAE considering AUC performances on contextual contamination and latent dimension of 32.
(a) Hidden layers
(b) Lambda
(c) Latent space
Figure 6: Study of RLAE hyperparameter: hidden layers number, lambda3 regulating local recovery term and latent space dimension.
Baseline
We propose to compare our approach against OCSVMSchölkopf et al. (2001), a simple one-class autoencoder (OC-AE), RSRAELai et al. (2020) and an ensemble autoencoder model RandAE based on Chen et al. (2017). For OCSVM OC-AE we use the implementation from PyODZhao et al. (2019b) and for RSRAE we use our PyTorch implementation of their code. All our results approximate theirs. We rigorously follow the guidelines provided by Lai et al. (2020).
Alternatively, we propose to use a variant of Algorithm 1 considering $P(y)=P(\zeta)$ for independent contamination. Finally we propose to benchmark our results on corpora presented in Table 1.
We also display CVDD Ruff et al. (2019b) results with several adaptations. In this approach, CVDD learn on all the training corpus instead of the training split. Considering this property, we use their implementation knowing the approach has advantage against other baseline models.
4.4 Results
We propose to present our results on three principal points: independent contamination, contextual contamination and robustness of model scores. Table 2 displays results on TAC1 (contamination of 0.1) and independent contamination. We can see that ErLA is outperforming all models. While the results are similar, our approach is notably standing over the others on SMS Spam and Reuters-21578. Thus, our ensemble method presents success for semantic related, spam and sentiment corpora.
Table 2 displays the experimental results conducted with our approach ErLA. We observe that our approach is outperforming others model with AUROC metric and AUPRC metric. We can see that usage of ErLA allow to mitigate unstable decision of the original RSRAE. We can also see significant difference of performance with Web of Science corpus and Reuters-21578. Additionally, we can observe that the original one-class autoencoder highly benefits from randomly connection and ensemble technique, as it close the gap with other models.
While our performances are competitive, the principal purpose of tackling outlier detection with ensemble methods is to mitigate the bias-variance trade-off. We propose to compare the model results with box-plots, similarly to the previous chapter. The main objective of our contribution is to robust outlier scores for contextual outliers with text. The Figure 2 and the Figure 3 displays an outperforming results from our approach. We can see that the variance of our model is noticeable as the box variance are always smaller than its competitors. Also, the min and max possible scores are close from the median scores, concluding to see that our approach is more efficient, more robust and can handle well language model like RoBERTA.
We can note that reference results are considerably different from ours. The principal reason is due the the contamination process for performing experimental study. We can observe that the literature contaminate both independent anomalies and contextual anomalies without distinctive analysis.
5 Discussion
In this section we propose an observation of various properties of our autoencoder RLAE and its impact inside ErLA (ensemble).
Base detector number
The Figure 4 displays the performance impact against the base detector number in ErLA. We can see that starting from 30, there is not substantial gains to increase the number of base detector in the ensemble. While it can slowly increase AUC, it considerably increase the computation time. 20 is then a concrete spot.
Number of local neighbours
In Figure 5 we display the impact of the hyperparameter $k$ in RLAE. As stated in the Reuters-21578 results, we can see that this hyperparameter needs to be wisely picked. While a low value seems to be preferable, we can see that our autoencoder is still robust against this hyperparameter.
Parameters
In Figure 6 is presented sensibility to three hyperparemeter that impact the learning step: number of hidden layer, the $\lambda_{3}$ parameter and the size of latent space in the ErLA model. While variance is tackled in ensemble methods, hidden layers are not a highly sensible parameter considering the computation/performance tradeoff. In Section 4 the hyperparameter $\lambda_{3}<\lambda_{1}$ is set to $0.05$. The curve is the AUC performance against $\lambda_{3}$ for $\lambda_{1}=0.1$. Finally, the learned embedding inside RLAE should be preferably lower than the output dimension of the autoencoder.
6 Conclusion
In this work we have introduced ErLA, an ensemble approach with robust autoencoders, optimized through LNE for tackling contextual anomaly in text. Conducted experiments demonstrate that our approach outperform all baseline method with less variance.
One perspective is to study the integration of attention head for mitigating the black box problem of our model. It is common, recently, to display text with their corresponding temperature, thanks to recent language model based on transformers.
The representation of text is a key concept that we want to investigate in the near future. Our approach has proven state of the art results and a great robustness against two kinds of outliers and with a small amount of available documents.
Furthermore, we have displayed that reference contributions have not put sufficient effort to the contamination process in their protocol. One promising perspective is to propose an unsupervised approach for generating different kinds of outliers. Also, our work mainly focuses on the semantic structure of text but syntax is also a promising direction.

### v2

@@
 Abstract
-In this work, a robust autoencoder ensemble-based approach designed to address anomaly detection in text corpora is introduced.
-Each autoencoder within the ensemble incorporates a local robust subspace recovery projection of the original data in its encoding embedding, leveraging the geometric properties of the k-nearest neighbors to optimize subspace recovery and identify anomalous patterns in textual data.
-The evaluation of such an approach needs an experimental setting dedicated to the context of textual anomaly detection.
-Thus, beforehand, a comprehensive real-world taxonomy is introduced to distinguish between independent anomalies and contextual anomalies.
-Such a study to identify clearly the kinds of anomalies appearing in a textual context aims at addressing a critical gap in the existing literature.
-Then, extensive experiments on classical text corpora have been conducted and their results are presented that highlights the efficiency, both in robustness and in performance, of the robust autoencoder ensemble-based approach when detecting both independent and contextual anomalies.
-Diverse range of tasks, including classification, sentiment analysis, and spam detection, across eight different corpora, have been studied in these experiments.
+Anomaly detection (AD) is a fast growing and popular domain among established applications like vision and time series.
+We observe a rich literature for these applications, but anomaly detection in text is only starting to blossom.
+Recently, self-supervised methods with self-attention mechanism have been the most popular choice.
+While recent works have proposed a working ground for building and benchmarking state of the art approaches, we propose two principal contributions in this paper: contextual anomaly contamination and a novel ensemble-based approach.
+Our method, Textual Anomaly Contamination (TAC), allows to contaminate inlier classes with either independent or contextual anomalies.
+In the literature, it appears that this distinction is not performed.
+For finding contextual anomalies, we propose RoSAE, a Robust Subspace Local Recovery Autoencoder Ensemble.
+All autoencoders of the ensemble present a different latent representation through local manifold learning.
+Benchmark shows that our approach outperforms recent works on both independent and contextual anomalies, while being more robust.
+We also provide 8 dataset comparison instead of only relying to Reuters and 20 Newsgroups corpora.
 A Robust Autoencoder Ensemble-Based Approach for Anomaly Detection in Text
@@
 1 Introduction
-Anomaly detection Chandola et al.
+Anomaly Detection (AD) Chandola et al.
 (2009); Ruff et al.
-(2021) and outlier detection Hawkins (1980); Hodge and Austin (2004); Zhang (2013); Aggarwal (2017a) are tasks aiming to estimate whether an observation is normal or not.
-Both of these tasks are referring to similar characteristics, and the terms anomaly and outlier are interchangeable in this paper.
-Several works have been devoted to study and categorise the different characteristics of an anomaly, and three principal kinds of anomalies can be noted: independent, contextual and collective.
-Defining a taxonomy for anomalies is crucial while the task to find anomalies lacks dedicated datasets.
-Such issue often leads practitioners to introduce an artificial contamination to produce anomalies into datasets.
-With the emergence of new machine learning methods and the availability of many datasets and corpora, anomaly detection can be addressed through many approaches Beggel et al.
-(2020); He et al.
-(2021); Goodge et al.
-(2023), however, in most cases, models are based on one-class classification (OCC) Khan and Madden (2014); Ruff et al.
-(2018); Sohn et al.
+(2021) and Outlier Detection (OD) Hodge and Austin (2004); Zhang (2013); Aggarwal (2017a) are tasks aiming to estimate whether an observation is normal or not.
+These tasks have been recently applied on text for highlighting particular or abnormal texts in a corpus.
+In this context, it is still a blossoming and promising field.
+Nowadays, most popular approaches are performing One-Class Classification (OCC) for discerning outliers (anomalies) from inliers Kannan et al.
+(2017); Lai et al.
+(2020); Fouché et al.
+(2020), and self-supervised approaches are often preferred for text Ruff et al.
+(2019); Manolache et al.
+(2021); Das et al.
+(2023).
+Self-supervised approaches are not limited to text, and applications with large language model for vision exists Wang et al.
+(2019); Zhou et al.
 (2023).
@@
 The detailed documents are the abstract of the news articles taken from sources like Reuters, New York times, BBC, … The first scenario is the apparition of a culture-related document in a medical feed, and the second scenario is a political document in the culture feed.
-Performing anomaly detection on textual data is less common than on many other types of data (images, time series and medical) but it comes with several useful applications that helps discerning wrong web content, hateful messages, spam or also errors in news feed.
-One of the main difficulties in this domain is the reproduction of experimental protocols and results from the literature Bejan et al.
-(2023).
-Indeed, there is a great difference between tackling independent outliers and contextual outliers Mahapatra et al.
-(2012); Fouché et al.
-(2020); Bejan et al.
-(2023) using semantic in text.
-For the former, anomaly classifier needs to differentiate two kinds of documents that come from unrelated topics (eg.
-sports and computer) but for the latter, they needs to detect a topic contaminated with another "sibling" (see Fig.
-1 for examples of this difference).
-It appears that most of the recent works contaminate corpora without addressing the problem of which kind of anomaly/outlier is added in this process Manevitz and Yousef (2001); Kannan et al.
-(2017); Ruff et al.
-(2019a); Lai et al.
-(2020).
-In this paper, based on recent literature, we address the problem of taxonomy after defining two kinds of anomaly: independent and contextual.
-Those anomalies are dependent of any semantic hierarchy and we propose in the experiments section to detail furthermore how they are benchmarked.
-For tackling both kind of anomalies, we propose a novel ensemble-based approach built with a robust autoencoder.
-This paper is composed as follows.
-In Section 2, recent works are presented.
-In Section 3, our approach to build an Ensemble robust subspace local recovery encoder is introduced with a description of its properties.
-Several experiments are presented in Section 4.
-Finally, Section 6 concludes and presents some future work.
+Ensemble-based approaches are not present in the literature, despite their great success on various kind of data Chen et al.
+(2017); Briskilal and Subalalitha (2022).
+Whereas those approaches benefits of decision robustness and reduced variance.
+Most of recent approaches are using Multi-Head Self-Attention Ruff et al.
+(2019); Manolache et al.
+(2021); Das et al.
+(2023), learning a representation from pre-trained language models as BERT-based models Devlin et al.
+(2019); Liu et al.
+(2019); Reimers and Gurevych (2019).
+An underlying problem of AD lies in the identification of different kinds of anomalies.
+Recent approaches does not usually consider these differences when building a detection model.
+While independent, contextual and collective anomalies have been identified for other kind of data Chandola et al.
+(2009); Aggarwal (2017b), it is different for text.
+Usually text anomalies are only considered as documents from one topic appearing in the wrong topic subset (Figure 1).
+However anomalies can have multiple forms, and independent anomalies are easier to find than contextual anomalies.
+In this paper we introduce a novel ensemble-based approach that performs contextual anomaly detection.
+In comparison to state-of-the-art approaches, we focus first on the identification of two kinds of anomaly, independent and contextual, and introduce the Textual Anomaly Contamination (TAC), a novel experimental setup.
+Also, our main contribution lies in the introduction of a novel approach for detecting anomalies which is based on a robust autoencoder ensemble.
+Our autoencoder, Robust Subspace Local Recovery AutoEncoder (RLAE), faces contaminated corpora (trained on both inliers and anomalies) and discriminates inliers from anomalies under the assumption that one distribution is highly contaminated.
+Thus, our Robust Subspace Local Recovery Autoencoder Ensemble (RoSAE) represents inliers in a low-dimensional subspace and proposes different manifold projections for one observation.
 2 Related works
@@
 (2019) or RoBERTa Liu et al.
-(2019), and large language models such as Llama Touvron et al.
+(2019), and large language models such as Llama Dubey et al.
 (2024) and OPT Zhang et al.
 (2022) have shown promising characteristics for anomaly detection.
-However, only few methods of the literature propose their use Manolache et al.
-(2021); Ruff et al.
-(2019a).
+However, only few methods of the literature propose their use Ruff et al.
+(2019); Manolache et al.
+(2021); Das et al.
+(2023).
 Other methods like One-Class Support Vector Machine (OCSVM) Schölkopf et al.
@@
 (2017) rely on a TF-IDF representation.
-Surprinsinly, recent methods are not using outlier ensemble methods Aggarwal and Sathe (2015); Zhao et al.
+Surprisingly, recent methods are not using outlier ensemble methods Aggarwal and Sathe (2015); Zhao et al.
 (2019a); Zimek et al.
 (2014) for performing anomaly/outlier detection with text data.
-Their usage in both Mixture of Experts Du et al.
-(2022) and ensemble approach Briskilal and Subalalitha (2022) in various NLP tasks highlights that anomaly detection in text data misses important approaches.
 AutoEncoders (AE) have been widely used for anomaly/outlier detection with high-dimensional data Chen et al.
@@
 The Robust Subspace Recovery Lerman and Maunu (2018); Rahmani and Atia (2017) is a robust manifold learning technique that map inlier distributions in a subspace where anomalies/outliers are at the edge.
-An autoencoder based on such an approach Lai et al.
+Based on such an approach, an autoencoder Lai et al.
 (2020) encodes observation from the input space in a robust subspace.
-One problem of such autoencoder is the difficulty to address scenarios with features that present similar range values (problem of entanglement).
-This entanglement is especially more difficult with contextual anomalies where documents shares similar semantic properties.
+One problem of such autoencoder is the difficulty to address scenarios with features that present similar range values (problem of entanglement) Wang et al.
+(2022).
+Entanglement is especially more difficult with contextual anomalies where documents shares similar semantic properties.
 Such problem can be tackled using a local projection from the nearest neighbours in the latent representation space.
@@
 (2019) can help to mitigate it.
-In this paper we introduce an ensemble-based approach with randomly connected autoencoders.
-We also present a new autoencoder, robust subspace local recovery autoencoder (RLAE), that performs robust subspace representation with local neighbouring recovery.
-In this approach, called Ensemble robust subspace Local recovery Autoencoder (ErLA), a RLAE ensemble, it can be seen as an ensemble of several subspace that aims to find normal data with different manifolds.
-It is based on the hypothesis that one distribution is highly contaminated, and that inliers lies in a low-dimensional subspace.
-3 ErLA: Ensemble robust subspace Local recovery Autoencoder
-This section presents our approach ErLA and a description of its properties.
+Figure 2: Overview of our global system, including contamination setup and our proposed approach RoSAE.
+A large number of contributions have settled anomaly/outlier taxonomies Zhang (2013); Aggarwal (2017b); Ruff et al.
+(2019) and several types of anomalies have been proposed in the literature: Independent anomaly, Contextual anomaly and Collective anomaly.
+Consequently, various types of anomalies frequently coexist within one corpus.
+A topic is defined as the main subject of a document Bejan et al.
+(2023).
+Depending on the document, there may be multiple subtopics within a broader category (e.g., a sports topic that encompasses football and tennis).
+Thus, accounting for this hierarchical structure, a form of contextual anomaly arises Fouché et al.
+(2020).
+Three different approaches have been used for performing corpus contamination.
+The first is a manually picked contamination Manevitz and Yousef (2001); Kannan et al.
+(2017), using text classification corpora like 20 Newsgroups, where topics are unrelated.
+A more practical contamination process is performed by recent works like Lai et al.
+(2020); Fouché et al.
+(2020), in which the inlier topic is selected and all other topics are considered anomalies.
+In this scenario, the hierarchical structure of topics is not handled, and related topics are mixed with each others.
+Finally, more recently, corpora have been prepared with careful consideration of topic hierarchy Ruff et al.
+(2019); Manolache et al.
+(2021); Das et al.
+(2023).
+We also note that there exists different kind of training procedure.
+In Ruff et al.
+(2019); Manolache et al.
+(2021) the proposed model is trained on inlier class, and then tested on contaminated split.
+For Lai et al.
+(2020); Fouché et al.
+(2020); Das et al.
+(2023), the training and testing step is performed on contaminated split.
+3 RoSAE: Robust Subspace Local Recovery Autoencoder Ensemble
 While robust subspace recovery autoencoders have successfully tackled anomaly detection in text corpora, they lack locality and geometry awareness for mitigating disentanglement and manifold collapse in transformer-based language models.
-Thus, we introduce RLAE which integrates locality in the latent representation through a local neighbouring method.
-The section is structured with a presentation of the randomly connected autoencoders, followed by a presentation of RSR loss.
-We then introduce the locally linear embedding loss term of ErLA before presenting its ensemble method.
-Finally, we present the representation of text.
+Thus, we introduce Robust Subspace Local Recovery AutoEncoder (RLAE) which integrates locality in the latent representation through a local neighbouring method before presenting our ensemble-based approach RoSAE.
 3.1 Randomly Connected One-Class Autoencoder
-Instead of using fully connected autoencoders, we propose to use randomly connected autoencoders.
-In the case of RSRAE, it is a novel approach and allow us to build ensemble autoencoders with different base detectors.
 Let $\mathrm{X}$ be a dataset of $\mathrm{N}$ instances such as $\mathrm{X}=\{\mathrm{x_{1}},...,\mathrm{x_{N}}\}$.
@@
 An Autoencoder is a neural networks in which the encoder $\mathcal{E}$ maps an instance $\mathrm{x_{i}}$ in a latent representation noted $\mathrm{z}_{i}=\mathcal{E}(\mathrm{x}_{i})\in\mathbb{R}^{e}$ of dimension $e$.
-The RSR layer is a linear transformation $\mathbf{A}\in\mathbb{R}^{d\times e}$ that reduces the dimension to $d$.
+The Robust Subspace Recovery (RSR) layer is a linear transformation $\mathbf{A}\in\mathbb{R}^{d\times e}$ that reduces the dimension to $d$.
 We denote $\mathrm{\hat{z}_{i}}$ the representation of $\mathrm{z_{i}}$ through the RSR layer, such as $\mathrm{\hat{z}_{i}}=\mathbf{A}\mathrm{z_{i}}\in\mathbb{R}^{d}$.
@@
 Similarly to Chen et al.
-(2017) we introduce autoencoders with random connection such as we increase the variance of our model.
-In the autoencoders ensemble each autoencoder has a random probability of having several of its connections to be cut.
+(2017) we introduce an autoencoder with random connection such as we increase the variance of our model.
+In the autoencoder ensembles, each autoencoder has a random probability of having several of its connections to be pruned.
 Thus, we set the probability of disconnection with a random rate in $[0.2,0.5]$.
 3.2 Robust Subspace Recovery Layer
-We present the RSR AutoEncoder that aims to robustly and nonlinearly reduce the dimension of the original data Lerman and Maunu (2018).
-The RSR layer maps the inliers around their original locations and the outliers far from their original locations.
+The RSR layer maps the inliers around their original locations, while outliers are projected far from their original locations Lerman and Maunu (2018); Lai et al.
+(2020).
 The loss function minimises the sum of the autoencoder loss function noted $L_{AE}$ with the RSR loss function noted $L_{RSR}$.
@@
 For performing the subspace recovery, we denote two terms that have different roles in the minimisation process.
-The first term enforces the RSR layer to be robust (PCA estimation) and the second enforces the projection to be orthogonal:
+Two terms are presented, with the first term enforcing the RSR layer to be robust (PCA estimation), and the second enforcing the projection to be orthogonal:
 $L^{q}_{RSR}(\mathbf{A})=$ $\lambda_{1}\sum^{N}_{i=1}{||\mathrm{z_{i}}-\mathbf{A}^{\mathrm{T}}\mathrm{\hat{z}_{i}}||^{q}_{2}}$
@@
 3.3 Robust Local Embedding
-Locally Linear Embedding (LLE) Roweis and Saul (2000); Chen and Liu (2011) is a popular dimensionality reduction technique that aims to preserve the local geometry of the data in a lower-dimensional subspace.
-It is based on the assumption that data points in a local neighbourhood can be linearly represented by their neighbouring data points.
-The LNE term in the loss function encourages the autoencoder to learn representations that preserve the relationships between data points in their local neighbourhoods.
-By doing so, it helps to project the Euclidean distance with its neighbours in the learned subspace.
-Based on Equation 2, the reconstruction loss function of RSRAE enforces robustness with $L^{1}_{AE}$ and the orthogonality with $L^{1}_{RSR}$.
+Locally Linear Embedding (LLE) Roweis and Saul (2000); Chen and Liu (2011) proposes the assumption that a local geometry can be linearly represented with surrounding observations.
+The Local Neighbouring Embedding (LNE) term in the loss function encourages the autoencoder to learn representations that preserve the relationships between data points and their local neighbours.
+Based on Equation 2, the reconstruction loss function enforces robustness with $L^{1}_{AE}$ and the orthogonality with $L^{1}_{RSR}$.
 Because the learned representation of the encoder is compressed in a $e$ dimension space, the locality of the subspace is not handled.
-For tackling this problem, we propose to introduce a third term to $L_{RLAE}$ based on a local neighbours embedding (LNE).
+Based on LNE, we introduce a third term.
 Given a set of data points $\{\mathrm{x_{i}}\}_{i=1}^{N}$ in the input space, the goal of LNE is to find a lower-dimensional representation $\{\mathrm{z_{i}}\}_{i=1}^{N}$ in the output space (the subspace learned by the autoencoder) such that the local relationships between data points are preserved.
@@
 The weights $w_{ij}$ can be computed using the least squares method to minimize the reconstruction error: $\min_{\mathbf{w_{i}}}\left\lVert\mathrm{x_{i}}-\sum_{j\in\mathcal{N}_{i}}w_{ij}\mathrm{x_{j}}\right\rVert_{2}^{2}$ subject to the constraint $\sum_{j\in\mathcal{N}_{i}}w_{ij}=1$.
-The LNE term encourages the autoencoder to find a representation for each data point as a combination of its k-nearest neighbours in the input space.
+The LNE term encourages the autoencoder to find a representation for each data point as a combination of its k-nearest neighbours in the latent space.
 With minimisation of the LNE term in the loss function, the autoencoder learns to preserve the local linear relationships, which ultimately helps to project the Euclidean distance with its neighbours.
@@
 The weight $w_{j}$ assigned to the neighbour $\mathrm{x_{ij}}$ in the local reconstruction of $\mathrm{x_{i}}$ is determined based on its corresponding distances.
-The inclusion of the LNE term in the loss function encourages the autoencoder to preserve the local geometric structure of the data in the learned subspace.
-Finally, the RLAE cost function is measured as follows:
+The inclusion of the LNE term in the loss function encourages the autoencoder to preserve the local geometric structure of the data in the latent subspace.
+Finally, the Robust Subspace Local Recovery AutoEncoder (RLAE) loss function is measured as follows:
 $L_{RLAE}(\mathcal{E},\mathbf{A},\mathcal{D})$ $=L_{RSRAE}(\mathcal{E},\mathbf{A},\mathcal{D})$
@@
 3.4 Ensemble Learning
-The main idea behind ensemble methods is that a combination of several models, also called base detectors, and their outputs is more robust than usage of a single model.
-Such robustness can be observed against the bias-variance tradeoff and also for tackling the issue of overfitting.
-Although the possibility to combine multiple base detectors is intuitive, the design of such approaches needs special attention regarding normalisation of outputs.
-In ErLA, we use the reconstruction error of each autoencoders and then we normalise each base detector scores through the standard deviation of one unit.
+The main idea behind ensemble methods is that a combination of several models, also called base detectors, is more robust than usage of a single model.
+In RoSAE, we use the reconstruction error of each autoencoders and then we normalise each base detector scores through the standard deviation of one unit.
 We then take the median value for each observation.
 3.5 Text Representation
-Recent works Ruff et al.
-(2019a); Manolache et al.
-(2021) have recorded their results on BERT-based language models.
-Thus, we use RoBERTa Liu et al.
-(2019) for text representation.
-The ErLA model is not based on the self-attention mechanism, such as for Ruff et al.
-(2019a); Manolache et al.
-(2021), and we use the implementation of Reimers and Gurevych (2019).
-4 Experiments
-In this section we present several experiments on both independent anomalies and contextual anomalies.
-We present all corpora and how we perform this study.
-The section is structured as follows: i) presentation of the experimental protocol for tackling independent and contextual anomalies; ii) presentation of all corpora and their principal characteristics; iii) complete description of experiment settings; iv) results of conducted benchmarks.
-Algorithm 1 TAC: Textual Anomaly Contamination
+Recent works have recorded their results on BERT-based language models.
+The RoSAE model is not based on the self-attention mechanism, such as for Ruff et al.
+(2019); Manolache et al.
+(2021), and we use the RoBERTa Sentence Transformer from Reimers and Gurevych (2019).
+Algorithm 1 TAC: Textual Anomaly Contamination (Independent)
 Inlier topic $\zeta$, corpus $\mathrm{X}$, split size $l$, contamination rate $\mathrm{\nu}$
@@
 while $|\mathrm{Z}|<c$ do
+$\triangleright$ Set = instead of $\neq$ for contextual
      if $\mbox{Parent}(\mathrm{y}_{i})\neq\mbox{Parent}(\zeta)\mbox{}$ then
@@
 return $\mbox{Shuffle}(\mathrm{Z})$
-4.1 Experimental protocol
-A large number of contributions have settled anomaly/outlier taxonomies Hawkins (1980); Hodge and Austin (2004); Zhang (2013); Aggarwal (2017b); Ruff et al.
-(2019a) several types of outliers have been proposed in the literature: Point anomaly, Conditional/Contextual anomaly and Collective/Group anomaly.
-A similar taxonomy can be applied to textual data.
-Consequently, various types of outliers frequently coexist within the documents of a given corpus.
-The definition of a topic can be assimilated to the subject matter that a document addresses.
-Depending on the document type, there may be multiple subtopics within a broader category (e.g., a sports topic that encompasses football and tennis).
-Thus, accounting for this hierarchical structure introduces a form of contextual anomaly.
-These contextual outliers may appear unremarkable in isolation but are considered outliers when associated with a small subset of the corpus.
-Collective anomaly poses challenges in terms of formalisation due to the contextual nature of textual data.
-To illustrate this, consider a legal document mentioning a football player, which would be anomalous if incorrectly appearing in a sports-related corpus.
-Point outliers represent observations that lack any meaningful relationship with other topics.
+4 Textual Anomaly Contamination (TAC)
+Text AD is lacking dedicated dataset, resulting in all state-of-the-art methods preparing benchmark corpora with synthetic contamination.
+We propose a contamination algorithm for performing contamination with independent and contextual anomalies (Figure 2).
+To illustrate this, let a legal document mentioning a football player, it would be anomalous if incorrectly appearing in a sports-related corpus.
+Independent anomalies represent observations that lack any meaningful relationship with other topics.
 Specifically, anomaly topics and inlier topics have different hierarchical parents within the category structure.
-Let a labelled document of a corpus $(\mathrm{x},y)\in\mathrm{X}\times\mathcal{Y}$ and $\zeta$ be the inlier category, and its corresponding subset $\mathrm{X}_{\zeta}\subseteq\mathrm{X}$.
+Let a corpus of text $\mathrm{X}$ with its labelled topic $\mathcal{Y}$, a labelled document of a corpus $(\mathrm{x},y)\in\mathrm{X}\times\mathcal{Y}$ and $\zeta$ the inlier category The corresponding inlier subset is noted $\mathrm{X}_{\zeta}\subseteq\mathrm{X}$.
 We define $\mathcal{A}$ the subset of all anomalies such as $\mathcal{A}\subset\mathrm{X}$.
@@
 First, an observation $\mathrm{x_{i}}$ is considered to be an anomaly if its parent topic is different of inlier parent topics such as:
-$\mathcal{A}_{p}(\zeta)=\{\mbox{P}(\zeta)\neq\mbox{P}(\mathrm{y})|(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y}\}$ (8)
+$\mathcal{A}_{p}(\zeta)=\{a|\mbox{P}(\zeta)\neq\mbox{P}(\mathrm{y}),(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y}\}$ (8)
 The second constraint corresponds to documents that do not lie in $\mathrm{X}_{\zeta}$ but share the same parent topic as $\zeta$.
@@
 We write:
-$\mathcal{A}_{c}(\zeta)=\{\mbox{P}(\zeta)=\mbox{P}(\mathrm{y})|(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y},\mathcal{A}\backslash\mathcal{A}_{p}\}$ (9)
-Dataset Topics Hierarchy
-20 Newsgroups 20 Yes
-DBpedia 14 14 Yes
-Reuters-21578 90 No
-Web of Science 134 Yes
-Enron 2 No
-SMS Spam 2 No
-IMDB 2 No
-SST2 2 No
-Table 1: Presentation of datasets from the literature on outlier detection and the inherent tasks.
-We describe these corpora by indicating the existence of a topic hierarchy in the labels of the original corpus.
+$\mathcal{A}_{c}(\zeta)=\{a|\mbox{P}(\zeta)=\mbox{P}(\mathrm{y}),(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y}\}$ (9)
+The Algorithm 1 presents TAC for independent anomalies and considers three hyperparameter: the contamination rate $\nu$, length of the contaminated split $l$ and the inlier topic noted $\zeta$.
+5 Experiments
+In this section we present several experiments on both independent anomalies and contextual anomalies.
+We present all corpora and how we perform this study.
+Results are commented and an ablation study is proposed.
 5.1 Data
-Although there are dedicated datasets for outlier detection, such as ODDS or UCI, they mainly provide multidimensional data, time series and computer vision data.
-Applications such as spam detection and text classification have a rich set of corpora available.
 Recent work Lai et al.
 (2020); Ruff et al.
-(2019a); Kannan et al.
+(2019); Kannan et al.
 (2017); Mahapatra et al.
 (2012) uses classification datasets such as Reuters-21578111 http://www.daviddlewis.com/resources/testcollections/reuters21578/ and 20 Newsgroups222 http://qwone.com/~jason/20Newsgroups/ with a dedicated preparation in order to compare their approaches.
-We use the corpus presented in the section and for each available category, we apply the preparation of independent outliers and contextual outliers with TAC (Algorithm 1).
-To be fair with each method and each dataset, we set the size of the preparation subset to $350$ and the results are averaged over $10$ of runs.
-The data is pre-processed by removing lower case and stop words.
-The train part of each corpus is used for training and evaluation.
-The TF-IDF model is applied to the whole corpus and only tokens that appear at least three times are kept in the vocabulary.
+For all corpora we apply data contamination of independent anomalies and contextual anomalies with TAC (Algorithm 1).
+To be fair with each method and each dataset, we set the size of the preparation subset to the maximum number of available document, and results are averaged over $10$ of runs.
+Data is pre-processed with lowercase and stop words removal.
+The train split of each corpus is used for training, and test split for evaluation.
 In a first step, we set $\nu=0.10$.
-Independent anomalies ($\nu=0.1$)
-Model 20New. Reuters WOS DBpedia Enron SMS Sp. IMDB SST2
-OCSVM 0.884 0.851 0.876 0.962 0.687 0.683 0.539 0.575
-OC-AE 0.630 0.721 0.655 0.599 0.484 0.683 0.517 0.499
-RandAE 0.622 0.689 0.636 0.575 0.466 0.687 0.503 0.486
-RSRAE 0.669 0.768 0.731 0.634 0.518 0.642 0.540 0.577
-CVDD 0.781 0.895 0.948 0.958 0.629 0.717 0.513 0.562
-RLEA 0.792 0.829 0.843 0.840 0.611 0.657 0.521 0.542
-ErLA 0.907 0.926 0.939 0.983 0.713 0.813 0.560 0.594
-Contextual anomalies ($\nu=0.1$)
-Model 20New. Reuters WOS DBpedia
-AP AUC AP AUC AP AUC AP AUC
-OCSVM 0.220 0.681 0.471 0.755 0.599 0.889 0.566 0.882
-OC-AE 0.175 0.592 0.351 0.683 0.162 0.577 0.334 0.714
-RandAE 0.164 0.579 0.337 0.662 0.147 0.544 0.319 0.694
-RSRAE 0.177 0.584 0.340 0.729 0.214 0.623 0.260 0.660
-CVDD 0.131 0.613 0.481 0.824 0.623 0.916 0.646 0.919
-RLEA 0.201 0.642 0.420 0.757 0.334 0.738 0.368 0.747
-ErLA 0.325 0.718 0.609 0.880 0.687 0.921 0.840 0.951
-Table 2: Results of state of the art models for independent and contextual anomalies with the contamination rate $\nu=0.10$.
-Area under ROC (AUC) is the default evaluation metric for independent contamination, and Average Precision (AP) is also added for contextual contamination.
-The experimental study is performed on Distill RoBERTA.
-Each result is performed on test split prepared through Algorithm 1.
 20 Newsgroups.
 For 20 Newsgroups we separate the subtopics into seven main topics: computer, forsale, motors, politics, religion, science, sports.
-We do not count the topic "forsale" for contextual outliers because it has no subtopics.
 Reuters-21578.
@@
 DBpedia 14.
-For DBpedia 14 we create the topic hierarchy based on the ontology provided333mappings.dbpedia.org/server/ontology/classes/ and has six parent topics.
+For DBpedia 14 Zhang et al. (2015) we create the topic hierarchy based on the provided ontology333mappings.dbpedia.org/server/ontology/classes/ and denote six parent topics.
 Web Of Science.
-Web of Science is often used as a reference for hierarchical classification and provides three levels of topic hierarchy.
-The third level topics are distributed among the corresponding first level parents.
-Thus, seven parent topics are present and for child topics that are associated with more than one parent, we keep the largest set of children and delete the others.
+Web of Science Kowsari et al.
+(2017) is often used as a reference for hierarchical classification and provides three levels of topic hierarchy.
+Thus, seven parent topics are present and for child topics that are associated with more than one parent, we only keep the largest set of children.
+Spam.
+We use Enron [^4] https://www.cs.cmu.edu/~enron/ dataset and SMS Spam Almeida et al. (2011) dataset. We prepare a train and test split before contaminating ham in spam, and then spam in ham.
+Sentiment.
+IMDB Maas et al.
+(2011) and SST2 Socher et al.
+(2013) are used for benchmarking independent contamination with more difficult scenarios.
+Preparation of both corpora is similarly performed than for Spam preparation.
+5.2 Preparation
+We use TAC for preparing contextual contamination on each candidate inliers with $\nu=0.1$.
+The maximum amount of available documents is retrieved $|D|$.
+All results are comapred against Area Under Curve ROC (AUC) and Average Precision (AP) metrics.
+We also keep the number of runs for each corpus and each split contamination to $10$.
+Baseline
+We propose to compare our approach against OCSVMSchölkopf et al.
+(2001), RSRAELai et al.
+(2020), an autoencoder ensemble model RandNet Chen et al.
+(2017), CVDD Ruff et al.
+(2019) and FATE Das et al.
+(2023).
+For OCSVM we use the implementation from PyODZhao et al.
+(2019b) and for RSRAE we use our PyTorch implementation of their code.
+All our results approximate theirs.
+We also display CVDD results with several adaptations.
+In the original paper, CVDD learns on all the training corpus (inlier only) instead of the training split only.
+Considering this property, we use their implementation knowing the approach has advantage against other baseline models.
+The same comment goes for FATE, which additionally use few-shot learning and benefits from anomaly awareness.
+Implementation
+For RoSAE and RandNet we set the number of base predictors to $20$.
+We provide the code of our approach [^5] https://anonymous.4open.science/r/RoSAE-COLING-B12E using the PyOD base implementation Zhao et al.
+(2019b).
+We propose an implementation based on the BaseDetector of PyOD so each of the compared model can be accordingly tested.
+We use the pruning methods from PyTorch for creating randomly connected autoencoders.
+Additionally, we also set hyperparameters $\lambda_{1}=0.1$, $\lambda_{2}=0.1$ and $\lambda_{3}=0.05$.
+For avoiding manifold collapse problem and degenerates solutions, we advise that $\lambda_{3}<\lambda_{1}$.
+We set the latent dimension of the LNE layer to $32$.
+On the other hand, we set the epoch number to $50$ and random connection probability between $[0.2,0.5]$.
+Independent anomalies ($\nu=0.1$)
+Model 20New. Reuters WOS DBpedia Enron SMS Sp. IMDB SST2
+OCSVM 0.884 0.851 0.876 0.962 0.687 0.683 0.539 0.575
+RandNet 0.622 0.689 0.636 0.575 0.466 0.687 0.503 0.486
+RSRAE 0.669 0.768 0.731 0.634 0.518 0.642 0.540 0.577
+CVDD 0.781 0.895 0.948 0.958 0.629 0.717 0.513 0.562
+FATE 0.911 0.926 0.918 0.951 0.484 0.683 0.517 0.529
+RLAE 0.792 0.829 0.843 0.840 0.611 0.657 0.521 0.542
+RoSAE 0.907 0.926 0.939 0.983 0.713 0.813 0.560 0.594
+Contextual anomalies ($\nu=0.1$)
+Model 20New. Reuters WOS DBpedia
+AP AUC AP AUC AP AUC AP AUC
+OCSVM 0.220 0.681 0.471 0.755 0.599 0.889 0.566 0.882
+RandNet 0.164 0.579 0.337 0.662 0.147 0.544 0.319 0.694
+RSRAE 0.177 0.584 0.340 0.729 0.214 0.623 0.260 0.660
+CVDD 0.131 0.613 0.481 0.824 0.623 0.916 0.646 0.919
+FATE 0.301 0.689 0.536 0.867 0.628 0.913 0.784 0.951
+RLAE 0.201 0.642 0.420 0.757 0.334 0.738 0.368 0.747
+RoSAE 0.325 0.718 0.609 0.880 0.687 0.921 0.840 0.951
+Table 1: Independent and contextual contamination results with contamination rate $\nu=0.10$.
+Area under ROC (AUC) is the default evaluation metric for independent contamination and Average Precision (AP) is also added for contextual contamination.
+Embedding is Distill RoBERTA and each result is performed on test split through Algorithm 1.
+FATE uses few-shot learning and has awareness of 5 anomalies.
 (a) DBpedia 14
@@
 (d) Reuters-21578
-Figure 2: Results of our experimental study with $\nu=0.1$, split size of $350$ and number of base detector of $25$. The performance metric is AUROC (AC) and the text representation is RoBERTA.
+Figure 3: Experimental study (AUC) with $\nu=0.05$, split size of $350$ and number of base detector of $25$.
 (a) DBpedia 14
@@
 (d) Reuters-21578
-Figure 3: Boxplots of results of our experimental study with $\nu=0.1$, split size of $350$ and number of base detector of $25$.
-The performance metric is AUPRC (AP) and the text representation is RoBERTA.
-4.3 Setup
-We use TAC for preparing contextual contamination on each candidate inliers possible with $\nu=0.1$ (TAC1) and a split size of min$(|D|,1000)$.
-All results are performed on AUROC and AUPRC reference works from the previous section.
-We integrate results of one-class autoencoder (OC-AE) and we also benchmark results on a randomly connected autoencoder ensemble (RandAE) Chen et al.
-(2017).
-The architecture is similar to Chen et al.
-(2017) and the autoencoders are following their settings.
-We also keep the number of runs for each corpus and each split contamination to $10$.
-Implementation
-For ErLA and RandAE we setup similarly than with the autoencoder and we set the number of base predictors to $20$.
-We provide the code of our approach [^4] https://github.com/jrmip/ErLA/ using the PyOD base implementation Zhao et al.
-(2019b).
-We propose an implementation based on the BaseDetector of PyOD so each of the compared model can be accordingly tested.
-Randomly connected neural networks can be difficult to implment.
-Thus, we have implemented three version of our randomly connected autoencoder: one with setting the prune connections to 0 before each backpropagation, another one with the usage of PyTorch built-in prune module and finally a version with Torch-Pruning Fang et al.
-(2023).
-Based on our experiments, we use the built-in Pytorch module for reproducibility.
-Additionally, we also set hyperparameters $\lambda_{1}=0.1$, $\lambda_{2}=0.1$ and $\lambda_{3}=0.05$.
-For avoiding manifold collapse problem and degenerates solutions, we advise that $\lambda_{3}<\lambda_{1}$.
-We set the latent dimension of the LNE layer to $32$.
-On the other hand, we set the epoch number to $50$ and random connection probability between $[0.2,0.5]$.
-(a) DBpedia 14
-(b) 20 Newsgroups
-(c) Web of Science
-(d) Reuters-21578
-Figure 4: Number of base detectors considering AUC performances on contextual contamination.
-(a) DBpedia 14
-(b) 20 Newsgroups
-(c) Web of Science
-(d) Reuters-21578
-Figure 5: Number of k-nearest neighbours for one RLAE considering AUC performances on contextual contamination and latent dimension of 32.
+Figure 4: Experimental study (AP) with $\nu=0.05$, split size of $350$ and number of base detector of $25$.
+5.3 Results
+We propose to present our results on three principal points: independent contamination, contextual contamination and robustness of model scores.
+Table 1 displays results on TAC and independent contamination.
+We can see that RoSAE succeeds to compete against recent contributions like FATE.
+While the results are similar, our approach is notably standing over the others on non topic-related corpora.
+Thus, our ensemble method presents success for semantic related, spam and sentiment corpora on independent contamination.
+For contextual contamination we observe that our approach is outperforming others model with AUC metric and AP metric.
+We can see that usage of RoSAE allows to mitigate unstable decision of the original RLAE.
+We can also see significant difference of performance with 20 Newsgroups corpus and Reuters-21578.
+For contextual contamination, our approach outperforms all models.
+On AP metric, RoSAE presents 9% higher results than top competitor, on average.
+(a) Number of base detector
+(b) Study of hyperparameter K
+Figure 5: Results of RoSAE with contextual contamination on size of ensemble and on different values for hyperparameter K.
 (a) Hidden layers
@@
 Figure 6: Study of RLAE hyperparameter: hidden layers number, lambda3 regulating local recovery term and latent space dimension.
-Baseline
-We propose to compare our approach against OCSVMSchölkopf et al.
-(2001), a simple one-class autoencoder (OC-AE), RSRAELai et al.
-(2020) and an ensemble autoencoder model RandAE based on Chen et al.
-(2017).
-For OCSVM OC-AE we use the implementation from PyODZhao et al.
-(2019b) and for RSRAE we use our PyTorch implementation of their code.
-All our results approximate theirs.
-We rigorously follow the guidelines provided by Lai et al.
-(2020).
-Alternatively, we propose to use a variant of Algorithm 1 considering $P(y)=P(\zeta)$ for independent contamination. Finally we propose to benchmark our results on corpora presented in Table 1.
-We also display CVDD Ruff et al.
-(2019b) results with several adaptations.
-In this approach, CVDD learn on all the training corpus instead of the training split.
-Considering this property, we use their implementation knowing the approach has advantage against other baseline models.
-4.4 Results
-We propose to present our results on three principal points: independent contamination, contextual contamination and robustness of model scores.
-Table 2 displays results on TAC1 (contamination of 0.1) and independent contamination.
-We can see that ErLA is outperforming all models.
-While the results are similar, our approach is notably standing over the others on SMS Spam and Reuters-21578.
-Thus, our ensemble method presents success for semantic related, spam and sentiment corpora.
-Table 2 displays the experimental results conducted with our approach ErLA.
-We observe that our approach is outperforming others model with AUROC metric and AUPRC metric.
-We can see that usage of ErLA allow to mitigate unstable decision of the original RSRAE.
-We can also see significant difference of performance with Web of Science corpus and Reuters-21578.
-Additionally, we can observe that the original one-class autoencoder highly benefits from randomly connection and ensemble technique, as it close the gap with other models.
-While our performances are competitive, the principal purpose of tackling outlier detection with ensemble methods is to mitigate the bias-variance trade-off.
-We propose to compare the model results with box-plots, similarly to the previous chapter.
-The main objective of our contribution is to robust outlier scores for contextual outliers with text.
-The Figure 2 and the Figure 3 displays an outperforming results from our approach.
-We can see that the variance of our model is noticeable as the box variance are always smaller than its competitors.
-Also, the min and max possible scores are close from the median scores, concluding to see that our approach is more efficient, more robust and can handle well language model like RoBERTA.
-We can note that reference results are considerably different from ours.
+While our performances are competitive, the principal purpose of tackling anomaly detection with ensemble methods is to mitigate the bias-variance trade-off.
+For benchmarking the robustness of each method, we use a limited amount of documents in training and a low contamination rate ($\nu=0.05$), presenting a difficult scenario to handle.
+The Figure 3 and the Figure 4 displays outperforming results from our approach.
+We can see that the variance of our model is noticeable as the box variance are always similar or smaller than its competitors.
+Also, the min and max possible scores are close from the median scores, concluding to see that our approach is more efficient and more robust.
+Reference results are considerably different from ours.
 The principal reason is due the the contamination process for performing experimental study.
 We can observe that the literature contaminate both independent anomalies and contextual anomalies without distinctive analysis.
-5 Discussion
-In this section we propose an observation of various properties of our autoencoder RLAE and its impact inside ErLA (ensemble).
+Another observation lies in the competitive results from FATE thanks to its few-shot learning mechanism.
+6 Ablation study
+In this section we propose an observation of various properties of our autoencoder RLAE and its impact inside RoSAE (ensemble).
 Base detector number
-The Figure 4 displays the performance impact against the base detector number in ErLA.
+The Figure 5(a) displays the performance impact against the base detector number in RoSAE.
 We can see that starting from 30, there is not substantial gains to increase the number of base detector in the ensemble.
@@
 Number of local neighbours
-In Figure 5 we display the impact of the hyperparameter $k$ in RLAE.
+In Figure 5(b) we display the impact of the hyperparameter $k$ in RLAE.
 As stated in the Reuters-21578 results, we can see that this hyperparameter needs to be wisely picked.
-While a low value seems to be preferable, we can see that our autoencoder is still robust against this hyperparameter.
+While a low value seems to be preferable, we can see that our autoencoder is still robust against this hyperparameter for other dataset.
 Parameters
-In Figure 6 is presented sensibility to three hyperparemeter that impact the learning step: number of hidden layer, the $\lambda_{3}$ parameter and the size of latent space in the ErLA model.
+In Figure 6 is presented sensibility to three hyperparemeter that impact the learning step: number of hidden layer, the $\lambda_{3}$ parameter and the size of latent space in the RoSAE model.
 While variance is tackled in ensemble methods, hidden layers are not a highly sensible parameter considering the computation/performance tradeoff.
@@
 7 Conclusion
-In this work we have introduced ErLA, an ensemble approach with robust autoencoders, optimized through LNE for tackling contextual anomaly in text.
-Conducted experiments demonstrate that our approach outperform all baseline method with less variance.
-One perspective is to study the integration of attention head for mitigating the black box problem of our model.
-It is common, recently, to display text with their corresponding temperature, thanks to recent language model based on transformers.
-The representation of text is a key concept that we want to investigate in the near future.
-Our approach has proven state of the art results and a great robustness against two kinds of outliers and with a small amount of available documents.
-Furthermore, we have displayed that reference contributions have not put sufficient effort to the contamination process in their protocol.
-One promising perspective is to propose an unsupervised approach for generating different kinds of outliers.
-Also, our work mainly focuses on the semantic structure of text but syntax is also a promising direction.
+In this work we have introduced RoSAE, an ensemble approach with robust autoencoders, optimized through LNE for tackling contextual anomaly in text.
+Conducted experiments demonstrate that our approach outperform all baseline methods with less variance on contextual contamination.
+We can also note that results displays that state-of-the-art approaches are performing independent contamination, and present less robust results than RoSAE.
+8 Limitations and Ethical Considerations
+We admit one major limitation from this work.
+As we have presented a large number of dataset, we lack of dedicated corpora.
+While our work is an attempt to reproduce real-world scenario, we only rely to synthetic contamination.
+On the other hand, anomaly detection is useful for users and experts, but can lead to unfortunate usages in social media.